\int \frac{2}{x} dx
1
Expert's answer
2020-09-17T15:31:16-0400

Solution. Find the integral


"\int \frac{2}{x} dx=2ln \vert x \rvert +C"

where C is constant.

Answer.

"\int \frac{2}{x} dx=2ln \vert x \rvert +C"

where C is constant.